# Tokenizing Buildings: A Transformer for Layout Synthesis

### v1

@@ 5 Conclusion
We introduced the SBM, a Transformer architecture tailored to BIM-like room layouts through BIM-Token Bundles, sparse attribute–feature matrices, and a mixed-type embedding module shared by encoder and decoder. This representation treats topology, walls, openings, and entities as sequence-native tokens with joint categorical and continuous attributes, enabling a single backbone to support both encoder-only room embeddings and an encoder–decoder pipeline (DDEP). On a BIM-grounded benchmark with specification-aware coverage, navigability, and overlap/clearance metrics, DDEP delivers near-complete inventory satisfaction, state-of-the-art navigability, and the lowest violation rates across LLM/VLM and domain-specific baselines. Encoder ablations show that SBM yields geometry- and entity-aware embeddings that complement generic text encoders, trading some raw ranking strength for substantially better within-type structure. Our study suggests that modestly sized, domain-specific sequence models over well-designed BIM tokenizations can outperform much larger general-purpose systems on layout quality.
@@
Advanced vertical constraints like sloped ceilings or stairs/elevation changes, and MEP/service routing are not modeled. The benchmark largely comprises residential rooms from our corpus, so distributional shifts (styles, scales, local codes) may affect external validity.

### v2

@@ 5 Conclusion
We introduced SBM, a Transformer that encodes BIM room layouts as sparse attribute–feature matrices of mixed categorical and continuous tokens, supporting both encoder-only room embeddings and encoder–decoder generation (DDEP) within a single backbone. DDEP achieves near-complete inventory satisfaction, state-of-the-art navigability, and the lowest violation rates across LLM/VLM and domain-specific baselines; encoder-only embeddings yield substantially better within-type clustering than generic text encoders. A controlled same-data benchmark confirms that these gains are architectural, with ablations identifying joint continuous-feature embedding and entity ordering as the primary drivers. These results suggest that modestly sized, domain-specific sequence models over well-designed BIM tokenizations can outperform much larger general-purpose systems. Because DDEP outputs are BIM-native tokens, they can be consumed by downstream agentic systems for iterative LLM-driven refinement, opening a path toward human-in-the-loop layout co-design.
@@
Advanced vertical constraints like sloped ceilings or stairs/elevation changes, and MEP/service routing are not modeled. Our controlled baselines (ATISS, BLT) require representational bridges from their native formats to our BIM tokenization, which may not fully preserve their design strengths. The benchmark comprises residential rooms from our corpus; existing public indoor-scene datasets (e.g., 3D-FRONT) use axis-aligned bounding boxes in global coordinates and lack the wall-referenced, parametric BIM entities our tokenization targets, precluding direct evaluation on those benchmarks. To enable reproducibility and follow-on work, we will release an anonymized dataset subset, evaluation suite, and code upon acceptance. Extending to non-residential typologies, diverse architectural styles, and varied local codes is left for future work.